Time Run # Target 15:46 01563 Pol. 3He for testing 16:47 20247 BeO for Optics 17:09 20248 Reference cell mistakenly filled with H2 17:19 20249 H2 cell still in 17:32 01564 H2 cell still in 17:33 01565 H2 cell still in 17:36 01566 H2 cell still in 17:39 01567 H2 cell still in 17:40 01568 H2 cell still in 17:42 01569 H2 cell still in 17:44 20250 H2 cell still in 17:44 01570 H2 cell still in 17:46 01571 H2 cell still in 17:47 01572 H2 cell still in 17:47 20251 H2 cell still in 17:49 01573 H2 cell still in 18:45 01574 H2 cell still in 18:45 20252 H2 cell still in 19:22 20253 H2 cell still in 19:22 01575 H2 cell still in 19:44 20254 H2 cell still in 20:43 01576 N2 Reference Cell Installed (H2 removed) 20:43 20255 N2 reference Cell 21:27 20256 N2 Reference Cell 21:27 01577 N2 Reference Cell
From there on it continued as normal with an N2 reference cell.
The d2n Run Database has been updated to reflect this.
A copy of this log entry has been emailed to: rsubedi@jlab.org, vasulk@jlab.org, ellie@jlab.org